### v1

@@ 6 Numerical experiments
In this section, we conduct some preliminary numerical experiments to evaluate the performance of our proposed AMPDA. All the experiments are conducted in MATLAB R2022a on a desktop equipped with an Intel Core i5-9500 CPU (3.00 GHz) and 16GB of RAM.
@@
In our experiments, we compare the performance of the proposed AMPDA with the gradient descent flow algorithm (GDFA) [38, Algorithm 1] for solving problems (2) and (3). The parameters of AMPDA are set as follows. We set $\underline{\alpha}=10^{-4}$, $\overline{\alpha}=10^{4}$, $\sigma=10^{-5}$, and $\gamma=1/2$ throughout the experiments. Motivated by the Barzilai-Borwein spectral method [3], we update the $\widetilde{\alpha}_{k}$ in Step 1 of AMPDA as
@@
(70) $\widetilde{\alpha}_{k}:=\max\left\{\underline{\alpha},\min\left\{\overline{\alpha},\frac{\|\Delta x^{k}\|_{2}^{2}}{|\langle\Delta x^{k},\Delta h_{1}^{k}\rangle|}\right\}\right\},$
where $\Delta x^{k}=x^{k}-x^{k-1}$ and $\Delta h_{1}^{k}=\nabla h_{1}(x^{k})-\nabla h_{1}(x^{k-1})$ if $k\geq 1$ and $|\langle\Delta x^{k},\Delta h_{1}^{k}\rangle|\neq 0$; otherwise, we set $\widetilde{\alpha}_{k}=1$. Let $T\in\mathbb{R}^{m\times m}$ be the diagonal matrix of which the $j$-th ($j=1,2,...,m$) diagonal entry is set as $T_{j,j}=1$ if $(\mathcal{T}_{\mu}(b))_{j}\neq 0$, and $T_{j,j}=0$ otherwise. For the AMPDA, we pick $i_{\star}\in\arg\max\{|(b-Tb)^{\top}a_{i}|:i\in\{1,2,\cdots,n\}\}$ with $a_{i}$ denoting the $i$-th column of $A$, compute
(71) $\hat{\theta}:=\frac{(b-Tb)^{\top}a_{i_{\star}}}{\|a_{i_{\star}}\|_{2}^{2}-\|Ta_{i_{\star}}\|_{2}^{2}},$
and then choose the initial point $x^{0}$ as
@@
(72) $x_{i}^{0}=\begin{cases}\arg\min\left\{|\theta-\hat{\theta}|:\theta\in[\underline{x}_{i_{\star}},\overline{x}_{i_{\star}}]\right\},&\text{ if }i=i_{\star},\\ 0,&\text{ else. }\end{cases}$
@@
We show in Appendix B that the $x^{0}$ given by (72) satisfies the condition (41) for both problem (2) and (3), if $A\in\mathbb{R}^{m\times n}~{}(m<n)$ has full row-rank, $b\notin\mathcal{S}_{\mu}$, and $\underline{x}<0<\overline{x}$. For the GDFA, the parameters are set as $\rho=2$, $\beta=10^{-6}$, and $\mathrm{jMax}=5$, following the notations in [38, Algorithm 1]. Also, the GDFA is initialized at the $x^{0}$ given by (72). All the tested algorithms are terminated when
$\frac{\|x^{k}-x^{k-1}\|_{2}}{\|x^{k}\|_{2}}<10^{-6}.$
Following similar settings to [43, Section 7.1], we generate a sensing matrix $A\in\mathbb{R}^{m\times n}$ with i.i.d. standard Gaussian entries and then normalize each column of $A$. Next, a subset of size $K^{\dagger}$ is chosen uniformly at random from $\{1,2,\cdots,n\}$ and a $K^{\dagger}$-sparse vector $x^{\dagger}\in\mathbb{R}^{n}$ with i.i.d. standard Gaussian entries on this subset is generated as the original true signal. The lower bound $\underline{x}$ and the upper bound $\overline{x}$ are set to be $-\max\{5,\|x^{\dagger}\|_{\infty}\}\mathbf{1}_{n}$ and $\max\{5,\|x^{\dagger}\|_{\infty}\}\mathbf{1}_{n}$, respectively, where $\mathbf{1}_{n}$ denotes the $n$-dimensional vector with all entries being $1$. We also produce $\widetilde{z}\in\mathbb{R}^{m}$ as a $\mu^{\dagger}$-sparse vector with $\mu^{\dagger}$ i.i.d. standard Gaussian entries at random (uniformly chosen) positions and set $z=2\mathrm{sign}(\widetilde{z})$. Finally, the measurement $b\in\mathbb{R}^{m}$ is generated as $b=Ax^{\dagger}-z+0.01\epsilon$, where $\epsilon\in\mathbb{R}^{m}$ has i.i.d. standard Gaussian entries.
@@
In our numerical tests, we set $(m,n,K^{\dagger},\mu^{\dagger})=(1280R,365R,40R,5R)$ with $R\in\{1,2,3,4\}$. For each $R$, we generate 50 random instances as described above. The model parameter $\lambda$ is set to $5$ and $0.5$ for problems (2) and (3), respectively. We set $K=\lceil 1.3K^{\dagger}\rceil$ for problem (3), and $\mu=\lceil 1.3\mu^{\dagger}\rceil$ for both problems.
Table 1: Results for solving problem (2).
Algorithm Time Iter Obj RecErr
$R=1$ AMPDA 0.005 39 5.057 2.77E-02
GDFA 0.278 32 5.057 2.77E-02
$R=2$ AMPDA 0.049 49 7.255 2.28E-02
GDFA 1.156 36 7.255 2.28E-02
$R=3$ AMPDA 0.156 55 8.892 2.20E-02
GDFA 2.761 40 8.892 2.18E-02
$R=4$ AMPDA 0.322 64 10.332 2.11E-02
GDFA 5.250 45 10.332 2.10E-02
Table 2: Results for solving problem (3).
Algorithm Time Iter Obj RecErr
$R=1$ AMPDA 0.005 31 1.105 3.68E-02
GDFA 1.068 142 1.105 3.65E-02
$R=2$ AMPDA 0.041 42 1.014 1.72E-02
GDFA 5.307 180 1.014 1.61E-02
$R=3$ AMPDA 0.142 53 1.019 1.98E-02
GDFA 16.098 246 1.019 2.01E-02
$R=4$ AMPDA 0.363 75 1.023 2.09E-02
GDFA 35.859 321 1.023 2.13E-02
@@
The computational results are summarized in Table 1 and Table 2, where we report the computational time in seconds (Time), the number of iterations (Iter), the objective value (Obj), and the recovery error ($\text{RecErr}=\|\widehat{x}-x^{\dagger}\|_{2}/\|x^{\dagger}\|_{2}$ with $\widehat{x}$ being the solution returned by the algorithms) averaged over 50 random instances. One can observe that the objective values and recovery error obtained by the AMPDA are comparable to those of the GDFA, while the AMPDA significantly outperforms the GDFA in terms of computational time for all the problem sizes $R\in\{1,2,3,4\}$. These results demonstrate the efficiency of the proposed AMPDA for solving the scale-invariant sparse signal recovery models (2) and (3).
@@ Appendix A The proof of (58)
^$\widehat{\partial}\Phi(\bar{x})\supseteq\widehat{\partial}\psi(\bar{x})+\varphi_{l}(\bar{x})\partial\varphi_{c}(\bar{x}),$
@@
(73) ^$\widehat{\partial}\Phi(\bar{x})=\left\{v:\mathop{\lim\inf}_{\begin{subarray}{c}x\to\bar{x}\\ x\neq\bar{x}\\ x\in\mathrm{dom}(\Phi)\end{subarray}}\frac{\Phi(x)-\Phi(\bar{x})-\langle v,x-\bar{x}\rangle}{\|x-\bar{x}\|_{2}}\geq 0\right\}.$
@@
(75) $\lim_{\begin{subarray}{c}x\to\bar{x}\\ x\neq\bar{x}\\ x\in\mathrm{dom}(\Phi)\end{subarray}}\frac{|r(x)|}{\|x-\bar{x}\|_{2}}=0,$
@@
(76) ^$\widehat{\partial}\Phi(\bar{x})$$\supseteq\left\{v:\mathop{\lim\inf}_{\begin{subarray}{c}x\to\bar{x}\\ x\neq\bar{x}\\ x\in\mathrm{dom}(\varphi_{l})\cap\mathrm{dom}(\varphi_{r})\end{subarray}}\frac{\psi(x)-\psi(\bar{x})-\langle v-\varphi_{l}(x)\bar{y},x-\bar{x}\rangle}{\|x-\bar{x}\|_{2}}\geq 0\right\}$
=^$=\widehat{\partial}\psi(\bar{x})+\varphi_{l}(\bar{x})\bar{y},$
for each $\bar{y}\in\partial\varphi_{c}(\bar{x})$. Consequently, this proposition is derived from (76).
@@
(77) $Q(x,y,z,c)=c^{2}f_{\mathcal{C}}(x)\Big{(}g^{*}(y)+2/c-\langle x,y\rangle\Big{)}+h_{1}(x)+h_{2}^{*}(z)-\langle x,z\rangle,$
@@
(78) $q(x,y,z,c):=c^{2}f_{\mathcal{C}}(x)\left(g^{*}(y)+2/c-\langle x,y\rangle\right)$
@@
(79) ^$\widehat{\partial}Q(\bar{x},\bar{y},\bar{z},\bar{c})=\widehat{\partial}q(\bar{x},\bar{y},\bar{z},\bar{c})+\left[\begin{array}[]{c}\nabla h_{1}(\bar{x})-\bar{z}\\ 0_{n}\\ \widehat{\partial}h_{2}^{*}(\bar{z})-\bar{x}\\ 0\end{array}\right],$
since $h_{1}(x)-\langle x,z\rangle$ is differentiable and $h_{2}$ is separable from $q$. Besides, with the help of Proposition A.1, it follows from (78) that
(80) ^$\widehat{\partial}q(\bar{x},\bar{y},\bar{z},\bar{c})\supseteq\widehat{\partial}q_{2}(\bar{x},\bar{y},\bar{z},\bar{c})+\left[\begin{array}[]{c}0_{n}\\ \bar{c}^{2}f_{\mathcal{C}}(\bar{x})\partial g^{*}(\bar{y})\\ 0_{n}\\ 0\end{array}\right],$
@@
(81) $q_{2}(x,y,z,c):=(g^{*}(\bar{y})+2/\bar{c}-\langle\bar{x},\bar{y}\rangle)c^{2}f_{\mathcal{C}}(x)+\bar{c}^{2}f_{\mathcal{C}}(\bar{x})\left(2/c-\langle x,y\rangle\right),$
@@
(82) ^$\widehat{\partial}q_{2}(\bar{x},\bar{y},\bar{z},\bar{c})\supseteq\left[\begin{array}[]{c}(g^{*}(\bar{y})+2/\bar{c}-\langle\bar{x},\bar{y}\rangle)\bar{c}^{2}\widehat{\partial}f_{\mathcal{C}}(\bar{x})\\ 0_{n}\\ 0_{n}\\ 2(g^{*}(\bar{y})+2/\bar{c}-\langle\bar{x},\bar{y}\rangle)f_{\mathcal{C}}(\bar{x})\bar{c}\end{array}\right]+\bar{c}^{2}f_{\mathcal{C}}(\bar{x})\left[\begin{array}[]{c}-\bar{y}\\ -\bar{x}\\ 0_{n}\\ -2/\bar{c}^{2}\end{array}\right].$

### v2

@@ 6 Numerical experiments
In this section, we conduct some preliminary numerical experiments to evaluate the performance of our proposed AMPDA. All the experiments are conducted in MATLAB R2025a on a desktop equipped with an Intel Core i5-9500 CPU (3.00 GHz) and 16GB of RAM.
@@
Throughout the experiments, the parameters of the proposed AMPDA are set as follows: we set $\underline{\alpha}=10^{-4}$, $\overline{\alpha}=10^{4}$, $\sigma=10^{-5}$, and $\gamma=1/2$. we update the $\widetilde{\alpha}_{k}$ in Step 1 of AMPDA as
@@
$\widetilde{\alpha}_{k}:=\max\left\{\underline{\alpha},\min\left\{\overline{\alpha},\frac{\|\Delta x^{k}\|_{2}^{2}}{|\langle\Delta x^{k},\Delta h_{1}^{k}\rangle|}\right\}\right\},$ (75)
where $\Delta x^{k}=x^{k}-x^{k-1}$ and $\Delta h_{1}^{k}=\nabla h_{1}(x^{k})-\nabla h_{1}(x^{k-1})$ if $k\geq 1$ and $|\langle\Delta x^{k},\Delta h_{1}^{k}\rangle|\neq 0$; otherwise, we set $\widetilde{\alpha}_{k}=1$. The attempted stepsize in (75) is designed following the original idea of the Barzilai–Borwein (BB) spectral method Barzilai-Borwein:IJNA1998 . Stepsize strategies of this BB type are well known to deliver remarkable practical efficiency in numerous optimization algorithms; see, for example, Lu-Xiao:SIAM2017 ; Raydan:1997SIOPT ; Wright-Nowak-Figueiredo:IEEESP-09 .
For problems (2) and (3), the initial point of all the tested algorithms is chosen as follows. Let $T\in\mathbb{R}^{m\times m}$ be the diagonal matrix whose $j$-th (j=1,2,,mj=1,2,...,m) diagonal entry is set to $T_{j,j}=1$ if $(\mathcal{T}_{\mu}(b))_{j}\neq 0$, and $T_{j,j}=0$ otherwise. We select $i_{\star}\in\arg\max\{|(b-Tb)^{\top}a_{i}|:i\in\{1,2,\cdots,n\}\}$, where $a_{i}$ denotes the $i$-th column of $A$, compute
$\hat{\theta}:=\frac{(b-Tb)^{\top}a_{i_{\star}}}{\|a_{i_{\star}}\|_{2}^{2}-\|Ta_{i_{\star}}\|_{2}^{2}},$ (76)
and then define the initial point $x^{0}$ by
@@
$x_{i}^{0}=\begin{cases}\arg\min\left\{|\theta-\hat{\theta}|:\theta\in[\underline{x}_{i_{\star}},\overline{x}_{i_{\star}}]\right\},&\text{ if }i=i_{\star},\\ 0,&\text{ else. }\end{cases}$ (77)
@@
In Appendix B, we show that this choice of $x^{0}$ satisfies the condition (43) for both problems (2) and (3), provided that $A\in\mathbb{R}^{m\times n}~(m<n)$ has full row-rank, $b\notin\mathcal{S}_{\mu}$, and $\underline{x}<0<\overline{x}$. All the tested algorithms are terminated once
$\frac{\|x^{k}-x^{k-1}\|_{2}}{\|x^{k}\|_{2}}<10^{-6}.$ (78)
6.1 Numerical results on the $L_{1}/L_{2}$-regularized least squares problem
In this subsection, we applied AMPDA for solving the $L_{1}/L_{2}$-regularized least squares problem
$\min\left\{\frac{\|x\|_{1}}{\|x\|_{2}}+\frac{\lambda}{2}\|Ax-b\|_{2}^{2}:x\neq 0,~\underline{x}\leq x\leq\overline{x},~x\in\mathbb{R}^{n}\right\},$ (79)
which is a special case of problem (2) with $\mu=0$. The test instances $(A,b)$ are obtained from the datasets leukemia, colon-cancer, and mpg in the LIBSVM repository [^3] https://www.csie.ntu.edu.tw/\~cjlin/libsvmtools/datasets/. In particular, as suggested in Huang-Jia-Yu:2010NIPS and following the approach in (Li-Sun-Toh:2018SIOPT, , Section 4.1), we expand the original features of the dataset mpg by an order 7 polynomial. For problem (79), we fix $\lambda=1$ and set the box constraint as $\underline{x}=-10\cdot\mathbf{1}_{n}$ and $\overline{x}=10\cdot\mathbf{1}_{n}$, where $\mathbf{1}_{n}$ denotes the $n$-dimensional vector of all ones.
We compare AMPDA with the following two methods. The first method is the alternating direction method of multipliers (ADMM) (Wang-Tao-Nagy-Lou:2021SIAM-ImageScience, , Algorithm 3.2), which is tailored for the problem (79). Following the notation in Wang-Tao-Nagy-Lou:2021SIAM-ImageScience , we set $\rho_{1}=\rho_{2}=\beta=\rho$, $\overline{\epsilon}=10^{-6}$, and $\texttt{j\_max}=5$. For better performance, we set $\rho=2$ for leukemia, $\rho=50$ for mpg, and $\rho=10$ for colon-cancer. The second one is the gradient descent flow method (GDFA) (Wang-Aujol-Gilboa-Lou:2024IPI, , Algorithm 1). Using the same notation as in (Wang-Aujol-Gilboa-Lou:2024IPI, , Algorithm 1), we set $\beta=10^{-6}$ and $\texttt{jMax}=5$. To obtain satisfactory numerical performance, we set $\rho=5$ for leukemia, $\rho=50$ for mpg, and $\rho=10$ for colon-cancer. We remark that the parameters of the proposed AMPDA follow the settings described at the beginning of Section 6 and remain unchanged for all datasets.
Table 1 summarizes the computational results together with the dimensions of the matrices $A$ constructed from the three datasets. For each method, we report the objective value (Obj), the number of iterations (Iter), and the computational time in seconds (Time), where the computational time is decomposed into two parts: the startup phase (reported in parentheses) and the iterative phase. Unlike the proposed AMPDA, which starts almost instantaneously, both GDFA and ADMM require longer startup times. This is mainly because these two methods perform Cholesky factorizations of certain matrices before the iterations begin, in order to accelerate solving the linear systems during the iterative phase. From Table 1, we observe that all three algorithms eventually converge to solutions of similar objective value on all tested datasets, while AMPDA consistently requires the least computational time. Figure 1 plots the objective value versus the computational time during the early stage of the iterative phase. From these curves, we observe that AMPDA consistently exhibits the fastest convergence among all methods. These results clearly demonstrate the superior effectiveness of AMPDA for solving the $L_{1}/L_{2}$-regularized least squares problems considered in this subsection.
Table 1: Results for the $L_{1}/L_{2}$-regularized least squares problem (79).
Dataset Alg Obj Iter Time
leukemia AMPDA 2.375 2856 4.35 (0.00)
$m=38,n=7129$ ADMM 2.379 994 268.84 (4.70)
GDFA 2.375 1575 286.75 (1.66)
colon-cancer AMPDA 5.347 6817 1.99 (0.00)
$m=62,n=2000$ ADMM 5.345 2843 82.03 (0.27)
GDFA 5.345 3380 54.80 (0.12)
mpg AMPDA 1.101 299 0.36 (0.00)
$m=392,n=3432$ ADMM 1.103 116 8.87 (0.61)
GDFA 1.101 181 7.55 (0.16)
(a) leukemia
(b) colon-cancer
(c) mpg
Figure 1: Objective value versus the computational time during the early stage of the iterative phase (excluding the startup phase) for problem (79) on the three datasets.
6.2 Numerical results on the robust signal recovery problem
[…]
@@ 6.2 Numerical results on the robust signal recovery problem
In our numerical tests, we set $(n,m,K^{\dagger},\mu^{\dagger})=(1280R,365R,40R,5R)$ with $R\in\{1,2,3,4\}$. For each value of $R$, we generate 50 random instances as described above. The model parameter $\lambda$ is set to $5$ and $0.5$ for problems (2) and (3), respectively. We set $K=\lceil 1.3K^{\dagger}\rceil$ for problem (3), and $\mu=\lceil 1.3\mu^{\dagger}\rceil$ for both problems. To assess the performance of AMPDA, we also include a comparison with the gradient descent flow algorithm (GDFA) (Wang-Aujol-Gilboa-Lou:2024IPI, , Algorithm 1). Using the same notation as in (Wang-Aujol-Gilboa-Lou:2024IPI, , Algorithm 1), we set the GDFA parameters in this experiment to $\rho=2$, $\beta=10^{-6}$, and $\texttt{jMax}=5$.
Table 2: Results for solving the robust signal recovery problem (2).
Alg Time Iter Obj RecErr
$R=1$ AMPDA 0.007(0.012) 39(7) 5.087(2.28E-01) 2.73E-02(7.21E-03)
GDFA 0.338(0.130) 32(5) 5.087(2.29E-01) 2.73E-02(7.14E-03)
$R=2$ AMPDA 0.049(0.008) 50(6) 7.260(1.98E-01) 2.29E-02(3.54E-03)
GDFA 1.175(0.082) 36(3) 7.260(1.98E-01) 2.29E-02(3.51E-03)
$R=3$ AMPDA 0.169(0.025) 56(6) 8.901(2.03E-01) 2.17E-02(1.77E-03)
GDFA 2.823(0.158) 41(3) 8.901(2.03E-01) 2.16E-02(1.76E-03)
$R=4$ AMPDA 0.336(0.032) 64(6) 10.359(1.99E-01) 2.08E-02(1.47E-03)
GDFA 5.375(0.287) 45(3) 10.359(1.99E-01) 2.07E-02(1.49E-03)
Table 3: Results for solving the robust signal recovery problem (3).
Alg Time Iter Obj RecErr
$R=1$ AMPDA 0.004(0.001) 30(4) 1.008(9.20E-04) 1.77E-02(5.25E-03)
GDFA 1.098(0.108) 143(14) 1.008(1.09E-03) 1.78E-02(5.67E-03)
$R=2$ AMPDA 0.040(0.007) 42(5) 1.014(7.50E-04) 1.73E-02(1.81E-03)
GDFA 5.314(0.559) 179(16) 1.014(7.82E-04) 1.59E-02(1.96E-03)
$R=3$ AMPDA 0.153(0.020) 54(7) 1.019(1.13E-03) 1.97E-02(1.18E-03)
GDFA 16.403(2.106) 248(33) 1.019(1.00E-03) 1.98E-02(1.22E-03)
$R=4$ AMPDA 0.366(0.044) 73(9) 1.023(9.53E-04) 2.09E-02(1.10E-03)
GDFA 36.253(5.024) 320(48) 1.023(9.00E-04) 2.11E-02(1.07E-03)
@@
The computational results are summarized in Table 2 and Table 3. For each method, we report the average computational time in seconds (Time), the number of iterations (Iter), the objective value (Obj), and the recovery error $\text{RecErr}=\|\widehat{x}-x^{\dagger}\|_{2}/\|x^{\dagger}\|_{2}$, where $\widehat{x}$ denotes the solution returned by the algorithm. In addition, the standard deviation over the 50 random instances is reported in parentheses next to each corresponding average. From these results, one can observe that AMPDA achieves objective values and recovery errors comparable to those of GDFA, while requiring substantially less computational time across all problem sizes $R\in\{1,2,3,4\}$. Furthermore, the standard deviations indicate that the performance of AMPDA is consistently stable over all test instances. These results demonstrate the efficiency of the proposed AMPDA for solving the scale-invariant sparse signal recovery models (2) and (3).
@@ Appendix A The proof of (63)
$\widehat{\partial}\Phi(\bar{x})\supseteq\widehat{\partial}\psi(\bar{x})+\varphi_{l}(\bar{x})\partial\varphi_{c}(\bar{x}),$
@@
$\widehat{\partial}\Phi(\bar{x})=\left\{v:\mathop{\lim\inf}_{\begin{subarray}{c}x\to\bar{x}\\ x\neq\bar{x}\\ x\in\mathrm{dom}(\Phi)\end{subarray}}\frac{\Phi(x)-\Phi(\bar{x})-\langle v,x-\bar{x}\rangle}{\|x-\bar{x}\|_{2}}\geq 0\right\}.$ (80)
@@
$\lim_{\begin{subarray}{c}x\to\bar{x}\\ x\neq\bar{x}\\ x\in\mathrm{dom}(\Phi)\end{subarray}}\frac{|r(x)|}{\|x-\bar{x}\|_{2}}=0,$ (82)
@@
$\widehat{\partial}\Phi(\bar{x})$ $\supseteq\left\{v:\mathop{\lim\inf}_{\begin{subarray}{c}x\to\bar{x}\\ x\neq\bar{x}\\ x\in\mathrm{dom}(\varphi_{l})\cap\mathrm{dom}(\varphi_{r})\end{subarray}}\frac{\psi(x)-\psi(\bar{x})-\langle v-\varphi_{l}(x)\bar{y},x-\bar{x}\rangle}{\|x-\bar{x}\|_{2}}\geq 0\right\}$ (83)
$=\widehat{\partial}\psi(\bar{x})+\varphi_{l}(\bar{x})\bar{y},$
for each $\bar{y}\in\partial\varphi_{c}(\bar{x})$. Consequently, this proposition is derived from (83). $\square$
@@
$Q(x,y,z,c)=c^{2}f_{\mathcal{C}}(x)\Big(g^{*}(y)+2/c-\langle x,y\rangle\Big)+h_{1}(x)+h_{2}^{*}(z)-\langle x,z\rangle,$ (84)
@@
$q(x,y,z,c):=c^{2}f_{\mathcal{C}}(x)\left(g^{*}(y)+2/c-\langle x,y\rangle\right)$ (85)
@@
$\widehat{\partial}Q(\bar{x},\bar{y},\bar{z},\bar{c})=\widehat{\partial}q(\bar{x},\bar{y},\bar{z},\bar{c})+\left[\begin{array}[]{c}\nabla h_{1}(\bar{x})-\bar{z}\\ 0_{n}\\ \widehat{\partial}h_{2}^{*}(\bar{z})-\bar{x}\\ 0\end{array}\right],$ (86)
since $h_{1}(x)-\langle x,z\rangle$ is differentiable and $h_{2}^{*}$ is separable from $q$. With the help of Proposition A.1, it follows from (85) that
$\widehat{\partial}q(\bar{x},\bar{y},\bar{z},\bar{c})\supseteq\widehat{\partial}q_{2}(\bar{x},\bar{y},\bar{z},\bar{c})+\left[\begin{array}[]{c}0_{n}\\ \bar{c}^{2}f_{\mathcal{C}}(\bar{x})\partial g^{*}(\bar{y})\\ 0_{n}\\ 0\end{array}\right],$ (87)
@@
$q_{2}(x,y,z,c):=(g^{*}(\bar{y})+2/\bar{c}-\langle\bar{x},\bar{y}\rangle)c^{2}f_{\mathcal{C}}(x)+\bar{c}^{2}f_{\mathcal{C}}(\bar{x})\left(2/c-\langle x,y\rangle\right),$ (88)
@@
$\widehat{\partial}q_{2}(\bar{x},\bar{y},\bar{z},\bar{c})\supseteq\left[\begin{array}[]{c}(g^{*}(\bar{y})+2/\bar{c}-\langle\bar{x},\bar{y}\rangle)\bar{c}^{2}\widehat{\partial}f_{\mathcal{C}}(\bar{x})\\ 0_{n}\\ 0_{n}\\ 2(g^{*}(\bar{y})+2/\bar{c}-\langle\bar{x},\bar{y}\rangle)f_{\mathcal{C}}(\bar{x})\bar{c}\end{array}\right]+\bar{c}^{2}f_{\mathcal{C}}(\bar{x})\left[\begin{array}[]{c}-\bar{y}\\ -\bar{x}\\ 0_{n}\\ -2/\bar{c}^{2}\end{array}\right].$ (89)